OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (trojanpoem):

y = -3 x + 3

OpenStudy (trojanpoem):

Do you want the steps ?

OpenStudy (anonymous):

Oh heyy thanks!! And yes that would help :D

OpenStudy (trojanpoem):

get the slope of the line y = -3x +12 which is - ( -x) / y = -3 / 1 ( y + 3 x - 12 = 0) y- y(1)/ x- x(1) = the slope

OpenStudy (trojanpoem):

Put everything in its place y - 6 = -3x - 3 y = -3 x + 3

OpenStudy (trojanpoem):

Cya

OpenStudy (anonymous):

Thanks!